A civic organization includes 200 members, who have an average income of $58,000, with a standard deviation of $10,000. A simple random sample of n 5 30 members is selected to participate in the annual fund-raising drive. What is the probability that the average income of the fund-raising group will be at least $60,000?



1
Expert's answer
2021-07-13T09:57:16-0400

"n=30," so "\bar{X}" is approximately normally distributed by the central limit theorem with


"E[\bar{X}]=\mu=58000"

"\sigma_{\bar{X}}=\dfrac{\sigma}{\sqrt{n}}\sqrt{\dfrac{N-n}{N-1}}=\dfrac{10000}{\sqrt{30}}\sqrt{\dfrac{200-30}{200-1}}"

"\approx1687.4748"

"P(\bar{X}\geq60000)=1-P(\bar{X}<60000)"

"=1-P(Z<\dfrac{60000-58000}{1687.4748})"

"\approx1-P(Z<1.1852)\approx0.1180"